You are moving in a car at 90 km/h and you see someone at the side of the highway. You can use your horn to warn the person. The frequency of the horn is 440 Hz. and the air temperature is 0°C. Calculate the frequency of the horn the person hears as you approach and pass. (Ans: Approaching: 480 Hz, Passing: 409 Hz)


1
Expert's answer
2022-05-18T08:31:24-0400

The Doppler's effect says

"f=f_0\frac{c\pm v_r}{c\pm v_s}"

(a) approaching

"f=f_0\frac{c}{c-v_s}"

"f=440\frac{331}{331-25}=480\:\rm Hz"

(b) passing

"f=f_0\frac{c}{c+v_s}"

"f=440\frac{331}{331+25}=409\:\rm Hz"